OpenStudy (anonymous):

Caley has a bag that contains 9 blue marbles and 13 yellow marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Caley selects a blue marble and then a yellow marble? Round your answer to the nearest percent. P(blue and yellow) ≈ ______%

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first one is blue 9 out of 22 or \(\frac{9}{22}\)

OpenStudy (misty1212):

second one is yellow: there are 13 yellow and only 21 in the bag, since the first one is gone, so 13 out of 21 or \(\frac{13}{21}\)

OpenStudy (misty1212):

multiply and find \[\frac{9}{22}\times \frac{13}{21}\]

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=9%2F22*13%2F21 looks like a bit above \(25\%\)

OpenStudy (anonymous):

Round it to the nearest percent is that what it would come out to be?

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

Im sorry I am doing a test and I got to finish it and hopefully master it